OpenStudy (anonymous):

what is the solution set of x^2-8x+52=0

OpenStudy (anonymous):

It is a quadratic equation which can be solved by either factorisation or using a quadratic fomula

OpenStudy (anonymous):

ok can you help factorize then

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What are the numbers which you can multiply to get 52 and add or subtract to get -8?? First lets right the factors of 52

OpenStudy (anonymous):

7 8

OpenStudy (anonymous):

!!

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

For this one it is better to use a quadratic formula

OpenStudy (anonymous):

4-6i,4+6i

OpenStudy (anonymous):

gotit

OpenStudy (anonymous):

\[x=\frac{ -b+or-\sqrt{(b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

\[ax ^{2}+bx+c\]

OpenStudy (theviper):

\[x^2-8x-52=0\] \[x^2-13x+4x-52=0\] \[x(x-13)+4(x-13)=0\] \[(x-13)(x+4)=0\]